OpenStudy (dragonbreath):

How many solutions does the equation have? 3(4x-3)-7x=5x-9 A. one solution B. infinite solutions C. no solution

OpenStudy (texaschic101):

3(4x - 3) - 7x = 5x - 9 --- distribute through the parenthesis 12x - 9 - 7x = 5x - 9 --- combine like terms 5x - 9 = 5x - 9 same line...means infinite solutions

OpenStudy (dragonbreath):

so its B.

OpenStudy (texaschic101):

yep..its B

OpenStudy (dragonbreath):

Thank you can you help me with 2 more